OpenStudy (anonymous):

A parallelogram with sides of 6 and 10 has an area of 30 square root of 2. Find the measure of each angle of the parallelogram. small angle ______° large angle _____°

OpenStudy (anonymous):

small angle: arcsin(30sqrt(2)/10/6) = 45 degrees large angle :arccos(30sqrt(2)/10/6) = 45 degrees + 90 degrees = 135 degrees

OpenStudy (amistre64):

area of a parallelgram is base * height 6sin(t) * 10 = 30sqrt(2)... yeah